# On the quantitative lower bounds for solutions to the Boltzmann equation in nonconvex domains

### v3

@@ Abstract.
In this article, we deal with the continuous mild solutions to the Boltzmann equation in a bounded space domain, with either angular cutoff assumption or non-cutoff assumption. Without assuming that the space domain is convex, we construct a Maxwellian lower bound for the cutoff case, and a weaker-than-Maxwellian lower bound for the non-cutoff case, thus extending results of [1, 2] in which the domain was assumed to be convex.
@@ 1. Introduction
In this article, we consider the lower bound of the mild solutions to the Boltzmann equation on a space domain $\Omega\subset\mathbb{R}^{3}$ satisfying
@@
The Boltzmann equation under the consideration reads
(1.1) $\partial_{t}f+v\cdot\nabla f=Q(f,f),\ \forall(t,x,v)\in[0,T)\times\Omega\times\mathbb{R}^{3},$
@@
(1.4) $Q(h_{1},h_{2})(v):=\int_{\mathbb{R}^{3}\times\mathbb{S}^{2}}B(|v-v_{*}|,\cos{\theta})[h_{2}(v^{\prime})h_{1}(v^{\prime}_{*})-h_{2}(v)h_{1}(v_{*})]dv_{*}d\sigma,$
@@
The problems of quantifying the positivity of the solutions to (1.1)-(1.3) have attracted a lot of attention from many authors. Beyond interest in physics, the problem is also of great importance for the study of kinetic theory. For example, the exponential lower bound is crucial in the study of the behavior of the solution to the related Laudau equation in [4, 5]. In [6], the Maxwellian lower bound is applied to derive the uniqueness of the solution to the Boltzmann equation in $L_{v}^{1}L_{x}^{\infty}(1+|v|^{2+0})$. In 1933, Carleman [3] showed the existence of exponential lower bounds on the radially symmetric solution of the spatially homogeneous Boltzmann equation with angular cutoff hard potentials. Since then, many results have been derived. For example, in [12] Pulvirenti and Wennberg improved the result by proving the Maxwellian lower bound for non-radially symmetric solutions of the spatially homogeneous Boltzmann equation with angular cutoff hard potentials. In [11], Mouhot removed the homogeneous assumption and obtained the Maxwellian lower bound for the torus domain. This result was extended to the bounded convex domain with $C^{2}$ boundary by Briant in [1, 2]. We note that in the articles [1, 2, 11], they also derived the ”less than Maxwellian” lower bounds for the non-cutoff case. Later, this result was improved to Maxwellian lower bounds in the case of hard and moderately soft potentials with the assumption that $\Omega$ is a torus by Imbert, Mouhot, Silvestre in [9] and in the case where $\Omega$ is $\mathbb{R}^{3}$ by Henderson, Snelson, and Tarfulea in [8].
However, the problem of Maxwellian lower bounds on a non-convex space domain remains an open problem in both the cutoff and non-cutoff cases. In this article, we derive a Maxwellian lower bound on an open bounded connected but not necessarily convex domain with a $C^{2}$ boundary for the cutoff case. For the non-cutoff case, we also derive a weaker than Maxwellian lower bound.
@@
(1.9) $Q^{+}\left(\mathbf{1}_{B(v_{0},\delta)},\mathbf{1}_{B(v_{0},\delta)}\right)\geq C\xi^{\frac{1}{2}}\mathbf{1}_{B(v_{0},\delta\sqrt{2}(1-\xi))},$
@@
In this article, we focus on the geometry of a $C^{2}$ connected domain and show that for any two points of $\Omega$ there always exists a suitable zigzag between these two points, which allows the spreading of the lower bound from one point to the other point like in [1, 2]. One difficulty comes from the fact that those zigzags in the domain must remain sufficiently far from the boundary of $\Omega$. In fact, the shape structure affects the amount of gas molecules passing through the boundary. In [1, 2], the lower bounds depend on: (1) hydrodynamic quantities. (2) the modulus of continuity of $f$. In our result, the lower bounds also depend on $\operatorname{Conn}_{d}(\Omega)$, the maximum number of ”good” zigzag segments needed to connect any two points in the space domain. We put the definition of $\operatorname{Conn}_{d}(\Omega)$ in (2.37) and address the detailed analysis concerning the geometry in Section 2.
From now on, until Section 6, we assume that $\nu<0$ (cutoff case), where $\nu$ is introduced in (1.8). To classify the type of solutions to (1.1)–(1.3) considered in this article, we first introduce a decomposition of the collision operator $Q$:
(1.10) $\begin{split}&Q(h_{1},h_{2})(v)\\ &=\int_{\mathbb{R}^{3}\times\mathbb{S}^{2}}B(|v-v_{*}|,\cos{\theta})[h_{2}(v^{\prime})h_{1}(v^{\prime}_{*})-h_{2}(v)h_{1}(v_{*})]dv_{*}d\sigma\\ &=\int_{\mathbb{R}^{3}\times\mathbb{S}^{2}}B(|v-v_{*}|,\cos{\theta})[h_{2}(v^{\prime})h_{1}(v^{\prime}_{*})]dv_{*}d\sigma-h_{2}(v)\int_{\mathbb{R}^{3}\times\mathbb{S}^{2}}B(|v-v_{*}|,\cos{\theta})h_{1}(v_{*})dv_{*}d\sigma\\ &=:Q^{+}[h_{1},h_{2}](v)-h_{2}(v)L[h_{1}](v),\end{split}$
@@
We assume that the domain $\Omega\subset\mathbb{R}^{3}$ satisfies Assumption A and the collision kernel $B$ satisfies Assumption B with $\nu<0$, where $\nu$ is introduced in (1.8). Given a nonnegative continuous function $f_{0}$ on $\overline{\Omega}\times\mathbb{R}^{3}$, we call a nonnegative function $f$ defined on $0,T)\times(\overline{\Omega}\times\mathbb{R}^{3})$ with $|f(t,x,v)|\leq C(1+|v|)^{-r}$ for some constant $C>0$ and $r>3$ for any $0<t\leq T$, $(x,v)\in\overline{\Omega}\times\mathbb{R}^{3}$ a ”continuous mild” solution to (1.1)–(1.3) with initial data $f_{0}$ if $f$ is continuous on $\Gamma_{conti}$ and for $(t,x,v)\in[0,T)\times\Omega\times\mathbb{R}^{3}$ and the function $f(t,x,v)$ satisfies
(1.12) $\begin{split}f(t,x,v)&=f_{0}(X_{0,t}(x,v),v)\exp[-\int_{0}^{t}L[f(s,X_{s,t}(x,v),\cdot)](v)ds]\\ &+\int_{0}^{t}\exp\left(-\int_{s}^{t}L[f(s^{\prime},X_{s^{\prime},t}(x,v),\cdot)](v)ds^{\prime}\right)\\ &\hskip 28.45274ptQ^{+}[f(s,X_{s,t}(x,v),\cdot),f(s,X_{s,t}(x,v),\cdot)](v)ds,\end{split}$
when $t\leq t_{\partial}:=\sup\{t>0|x-vs\in\overline{\Omega},\ \forall s\in[0,t]\}$, and
(1.13) $\begin{split}f(t,x,v)&=\alpha f(t-t_{\partial},X_{t-t_{\partial},t}(x,v),R(X_{t-t_{\partial},t}(x,v),v))\\ &\hskip 14.22636pt\exp\left[-\int_{t-t_{\partial}}^{t}L[f(s,X_{s,t}(x,v),\cdot)](v)ds\right]\\ &+(1-\alpha)\left(\int_{w\cdot n(X_{t-t_{\partial},t}(x,v))>0}f(t,X_{t-t_{\partial},t}(x,v),w)(w\cdot n(X_{t-t_{\partial},t}(x,v)))dw\right)\\ &\hskip 51.21504pt\frac{1}{2\pi T_{B}^{2}}e^{-\frac{|v|^{2}}{2T_{B}}}\exp\left[-\int_{t-t_{\partial}}^{t}L[f(s,X_{s,t}(x,v),\cdot)](v)ds\right]\\ &+\int_{t-t_{\partial}}^{t}\exp\left(-\int_{s}^{t}L[f(s^{\prime},X_{s^{\prime},t}(x,v),\cdot)](v)ds^{\prime}\right)\\ &\hskip 39.83368ptQ^{+}[f(s,X_{s,t}(x,v),\cdot),f(s,X_{s,t}(x,v),\cdot)](v)ds,\end{split}$
when $t\geq t_{\partial}$. Here, a detailed definition of the characteristic line $X_{s,t}(x,v)$ is provided in the Appendix.
Remark 1.
In case when the specular reflection boundary condition holds, i.e., $\alpha=1$, we can also write the continuous mild solutions of (1.1)–(1.3) as the function, continuous on $\Gamma_{conti}$, $f$ such that for all $(t,x,v)\in[0,T]\times\overline{\Omega}\times\mathbb{R}^{3}$,
(1.14) $\begin{split}f(t,x,v)&=f_{0}(X_{0,t}(x,v),V_{0,t}(x,v))\exp[-\int_{0}^{t}L[f(s,X_{s,t}(x,v),\cdot)](V_{s,t}(x,v))ds]\\ &+\int_{0}^{t}\exp\left(-\int_{s}^{t}L[f(s^{\prime},X_{s^{\prime},t}(x,v),\cdot)](V_{s^{\prime},t}(x,v))ds^{\prime}\right)\\ &\hskip 28.45274ptQ^{+}[f(s,X_{s,t}(x,v),\cdot),f(s,X_{s,t}(x,v),\cdot)](V_{s,t}(x,v))ds,\end{split}$
@@
Remark 2.
It is worth mentioning that in [7], in the diffusive reflection boundary condition ($\alpha=0$), the existence of a solution of bounded variation is established. However, in [10], it is shown that, unlike the convex case, the solution for diffusive reflection boundary condition is not continuous, with the singularity propagating along the grazing set. Thus, unlike the continuity assumption of [1, 2], we consider the set $\Gamma^{0}_{-}$, which is inspired by [10]. It is worth mentioning that in the same article, the continuity of solutions is ensured in diffuse reflection boundary conditions, with continuity of initial data and addition conditions.
@@
Given $\Omega\subset\mathbb{R}^{3}$ satisfying Assumption A, kernel $B$ satisfying Assumption B with $\nu<0$, $\alpha\in[0,1]$, a non-negative continuous function $f_{0}$ on $\overline{\Omega}\times\mathbb{R}^{3}$, we assume a continuous mild solution of (1.1)–(1.3) on $0,T)\times\overline{\Omega}\times\mathbb{R}^{3}$, with $T>0$ (and initial condition $f_{0}$) which satisfies the following property:
@@
Then,the following lower bound property holds: For any $\tau\in(0,T)$ there exist $\rho>0$ and $\theta>0$ which depend on $M$, $C_{\Phi},c_{\Phi},\gamma,b_{0},\nu,E_{f}$ (and $L_{f,p}$ if $\gamma<0$ ), $\tau$, $\alpha$, $\Omega$ and the modulus of continuity of $f$, such that
@@
For the non-cutoff case, the result is stated in Chapter 6. We note that the numbers $\rho$ and $\theta$ are computable thanks to the special zigzag which will be introduced in Proposition 2.5, through which we can guarantee that a certain amount and range of the lower bound can be generated though the propagation in Proposition 3.4 and derive a computable uniform lower bound (which depends on $\Omega$ since the construction of zigzag is affected by the geometric properties of $\Omega$) as in Proposition 3.5.
The rest part of the article is organized as follows. In Section 2, we analyze the boundary of non-convex domain to derive some key properties. In Section 3, we focus on the derivation of a series of diluting initial lower bounds and the propagation effect of lower bounds. In Section 4 and Section 5, we provide a detailed proof of the Maxwellian lower bound on mild solution to the Boltzmann equation with non-fully specular ($0\leq\alpha<1$) and fully specular reflection boundary condition ($\alpha=1$), respectively. In Section 6, a weaker exponential lower bound of solutions to the Boltzmann equation for non cutoff case is provided.

### v4

@@ Abstract.
In this article, we study the continuous mild solutions to the Boltzmann equation in a bounded spatial domain, under either angular cutoff assumption or non-cutoff assumption. Without assuming convexity of the spatial domain, we establish a Maxwellian lower bound in the cutoff case, and a weaker-than-Maxwellian lower bound for the non-cutoff case. This extends the results of [1, 2], where the convexity of the domain was required.
@@ 1. Introduction
In this paper, we investigate the lower bounds of the mild solutions to the Boltzmann equation on a spatial domain $\Omega\subset\mathbb{R}^{3}$ satisfying
@@
The Boltzmann equation under consideration reads
(1.1) $\partial_{t}f(t,x,v)+v\cdot\nabla f(t,x,v)=Q[f,f](t,x,v),\ \forall(t,x,v)\in[0,T)\times\Omega\times\mathbb{R}^{3},$
@@
(1.4) $Q[h_{1},h_{2}](v):=\int_{\mathbb{R}^{3}\times\mathbb{S}^{2}}B(|v-v_{*}|,\cos{\theta})[h_{2}(v^{\prime})h_{1}(v^{\prime}_{*})-h_{2}(v)h_{1}(v_{*})]\,dv_{*}\,d\sigma,$
@@
The problems of quantifying the positivity of the solutions to (1.1)-(1.3) have attracted considerable attention from many authors. Beyond interest in physics, the problem also plays a crucial role in kinetic theory. For example, the exponential lower bound is crucial in the study of the behavior of the solution to the related Laudau equation in [4, 5]. In [6], the Maxwellian lower bound is applied to derive the uniqueness of the solution to the Boltzmann equation in $L_{v}^{1}L_{x}^{\infty}(1+|v|^{2+0})$. In 1933, Carleman [3] showed the existence of exponential lower bounds on the radially symmetric solution of the spatially homogeneous Boltzmann equation with angular cutoff hard potentials. Since then, many results have been derived. For example, in [12] Pulvirenti and Wennberg improved the result by proving the Maxwellian lower bound for non-radially symmetric solutions of the spatially homogeneous Boltzmann equation with angular cutoff hard potentials. In [11], Mouhot removed the homogeneous assumption and obtained the Maxwellian lower bound for the torus domain. This result was extended to the bounded convex domain with $C^{2}$ boundary by Briant in [1, 2]. We note that in the articles [1, 2, 11], they also derived the ”less than Maxwellian” lower bounds for the non-cutoff case. Later, this result was improved to Maxwellian lower bounds in the case of hard and moderately soft potentials with the assumption that $\Omega$ is a torus by Imbert, Mouhot, Silvestre in [9] and in the case where $\Omega$ is $\mathbb{R}^{3}$ by Henderson, Snelson, and Tarfulea in [8].
However, the problem of Maxwellian lower bounds on a non-convex spatial domain remains an open problem in both the cutoff and non-cutoff cases. In this article, we derive a Maxwellian lower bound on an open bounded connected but not necessarily convex domain with a $C^{2}$ boundary for the cutoff case. For the non-cutoff case, we also derive a weaker than Maxwellian lower bound.
@@
(1.9) $Q^{+}\left[\mathbf{1}_{B(v_{0},\delta)},\mathbf{1}_{B(v_{0},\delta)}\right](v)\geq C\xi^{\frac{1}{2}}\mathbf{1}_{B(v_{0},\delta\sqrt{2}(1-\xi))}(v),$
@@
In this article, we focus on the geometry of a $C^{2}$ connected domain and show that for any two points of $\Omega$ there always exists a suitable zigzag between these two points, which allows the spreading of the lower bound from one point to the other point like in [1, 2]. One major difficulty arises from the requirement that these zigzags in the domain must remain sufficiently distant from the boundary of $\Omega$. In fact, the shape structure affects the amount of gas molecules passing through the boundary. In [1, 2], the lower bounds depend on: (1) hydrodynamic quantities. (2) the modulus of continuity of $f_{0}$. In our result, the lower bounds also depend on $\operatorname{Conn}_{d}(\Omega)$, the maximum number of ”good” zigzag segments needed to connect any two points in the spatial domain. We put the definition of $\operatorname{Conn}_{d}(\Omega)$ in (2.37) and address the detailed analysis concerning the geometry in Section 2. Our main contribution lies in showing that the Maxwellian lower bound can be propagated throughout a non-convex domain by exploiting geometric connectivity properties, quantified through $\operatorname{Conn}_{d}(\Omega)$.
Throughout Sections 1–5, we assume that $\nu<0$ (cutoff case), where $\nu$ is defined in (1.8). To classify the type of solutions to (1.1)–(1.3), we first decompose the collision operator $Q$ as follows:
(1.10) $\begin{split}&Q[h_{1},h_{2}](v)\\ =&\int_{\mathbb{R}^{3}\times\mathbb{S}^{2}}B(|v-v_{*}|,\cos{\theta})[h_{2}(v^{\prime})h_{1}(v^{\prime}_{*})-h_{2}(v)h_{1}(v_{*})]\,dv_{*}\,d\sigma\\ =&\int_{\mathbb{R}^{3}\times\mathbb{S}^{2}}B(|v-v_{*}|,\cos{\theta})[h_{2}(v^{\prime})h_{1}(v^{\prime}_{*})]\,dv_{*}\,d\sigma\\ &-h_{2}(v)\int_{\mathbb{R}^{3}\times\mathbb{S}^{2}}B(|v-v_{*}|,\cos{\theta})h_{1}(v_{*})\,dv_{*}\,d\sigma\\ =&:Q^{+}[h_{1},h_{2}](v)-h_{2}(v)L[h_{1}](v),\end{split}$
@@
We assume that the domain $\Omega\subset\mathbb{R}^{3}$ satisfies Assumption A and the collision kernel $B$ satisfies Assumption B with $\nu<0$, where $\nu$ is introduced in (1.8). Given a nonnegative, continuous function $f_{0}$ on $\overline{\Omega}\times\mathbb{R}^{3}$, we call a nonnegative function $f$ defined on $0,T)\times(\overline{\Omega}\times\mathbb{R}^{3})$ with $|f(t,x,v)|\leq C(1+|v|)^{-r}$ for some constant $C>0$ and $r>3$ for any $0<t\leq T$, $(x,v)\in\overline{\Omega}\times\mathbb{R}^{3}$ a ”continuous mild” solution to (1.1)–(1.3) with initial data $f_{0}$ if $f$ is continuous on $\Gamma_{conti}$ and for $(t,x,v)\in[0,T)\times\Omega\times\mathbb{R}^{3}$ and the function $f(t,x,v)$ satisfies the following integral identities:
(1.12) $\begin{split}f(t,x,v)&=f_{0}(X_{0,t}(x,v),v)\exp\left(-\int_{0}^{t}L[f(s,X_{s,t}(x,v),\cdot)](v)\,ds\right)\\ &+\int_{0}^{t}\exp\left(-\int_{s}^{t}L[f(s^{\prime},X_{s^{\prime},t}(x,v),\cdot)](v)\,ds^{\prime}\right)\\ &\hskip 28.45274ptQ^{+}[f(s,X_{s,t}(x,v),\cdot),f(s,X_{s,t}(x,v),\cdot)](v)\,ds,\end{split}$
when $t\leq t_{\partial}(x,v):=\sup\{t\geq 0|x-vs\in\overline{\Omega},\ \forall s\in[0,t]\}$, and
(1.13) $\begin{split}f(t,x,v)&=\alpha f(t-t_{\partial}(x,v),X_{t-t_{\partial}(x,v),t}(x,v),R(X_{t-t_{\partial}(x,v),t}(x,v),v))\\ &\hskip 14.22636pt\exp\left(-\int_{t-t_{\partial}(x,v)}^{t}L[f(s,X_{s,t}(x,v),\cdot)](v)\,ds\right)\\ &+(1-\alpha)\left(\int_{w\cdot n(X_{t-t_{\partial}(x,v),t}(x,v))>0}f(t,X_{t-t_{\partial}(x,v),t}(x,v),w)(w\cdot n(X_{t-t_{\partial}(x,v),t}(x,v)))dw\right)\\ &\hskip 51.21504pt\frac{1}{2\pi T_{B}^{2}}e^{-\frac{|v|^{2}}{2T_{B}}}\exp\left(-\int_{t-t_{\partial}(x,v)}^{t}L[f(s,X_{s,t}(x,v),\cdot)](v)\,ds\right)\\ &+\int_{t-t_{\partial}(x,v)}^{t}\exp\left(-\int_{s}^{t}L[f(s^{\prime},X_{s^{\prime},t}(x,v),\cdot)](v)\,ds^{\prime}\right)\\ &\hskip 39.83368ptQ^{+}[f(s,X_{s,t}(x,v),\cdot),f(s,X_{s,t}(x,v),\cdot)](v)\,ds,\end{split}$
when $t\geq t_{\partial}(x,v)$. Here, a detailed definition of the characteristic line $X_{s,t}(x,v)$ is provided in the Appendix.
Remark 1.1.
In the case of specular reflection ($\alpha=1$), the continuous mild solution of (1.1)–(1.3) can also be written as the continuous function on $\Gamma_{conti}$, $f$ such that for all $(t,x,v)\in[0,T]\times\overline{\Omega}\times\mathbb{R}^{3}$,
(1.14) $\begin{split}f(t,x,v)&=f_{0}(X_{0,t}(x,v),V_{0,t}(x,v))\exp\left(-\int_{0}^{t}L[f(s,X_{s,t}(x,v),\cdot)](V_{s,t}(x,v))\,ds\right)\\ &+\int_{0}^{t}\exp\left(-\int_{s}^{t}L[f(s^{\prime},X_{s^{\prime},t}(x,v),\cdot)](V_{s^{\prime},t}(x,v))\,ds^{\prime}\right)\\ &\hskip 28.45274ptQ^{+}[f(s,X_{s,t}(x,v),\cdot),f(s,X_{s,t}(x,v),\cdot)](V_{s,t}(x,v))\,ds,\end{split}$
@@
Remark 1.2.
The continuity assumption of the solution $f(t,x,v)$ is technical, as the proof only involves the continuity of the initial condition. Nonetheless, in the diffusive reflection boundary condition ($\alpha=0$), the existence of a solution of bounded variation is established in [7]. However, as shown in [10], unlike in the convex case, the solution loses continuity, with singularities propagating along the grazing set. Motivated by [10], we therefore consider the set $\Gamma_{conti}$. In the same work, the authors also showed that the continuity of solutions on $\Gamma_{conti}$ is ensured under diffuse reflection boundary conditions, with continuity of the initial data and addition conditions.
@@
Suppose that $\Omega\subset\mathbb{R}^{3}$ satisfies Assumption A and that the kernel $B$ satisfies Assumption B with $\nu<0$. We consider a non-negative continuous function $f_{0}$ on $\overline{\Omega}\times\mathbb{R}^{3}$. Let $f(t,x,v)$ be a continuous mild solution of (1.1)–(1.3) on $0,T)\times\overline{\Omega}\times\mathbb{R}^{3}$, with initial condition $f_{0}$ for some $T>0$, and let $\alpha\in[0,1]$. We assume that $f(t,x,v)$ satisfies the following properties:
@@
Then, the following lower bound holds: There exists $0<\tau_{0}\leq T$ such that for any $\tau\in(0,\tau_{0})$, there exist $\Delta_{\tau_{0}}>0$, $\rho>0$ and $\theta>0$ depending on $\tau_{0}$, $M$, $C_{\Phi},c_{\Phi},\gamma,b_{0},\nu,E_{f}$ (and $L_{f,p}$ if $\gamma<0$ ), $\tau$, $\alpha$, $\Omega$ and on the modulus of continuity of $f_{0}$, such that
(1.19) $f(t,x,v)\geq\frac{\rho}{(2\pi\theta)^{\frac{3}{2}}}e^{-\frac{|v|^{2}}{2\theta}},\ \forall\,t\in[\tau,\Delta_{\tau_{0}}),\ \forall\,(x,v)\in\overline{\Omega}\times\mathbb{R}^{3}.$
The next theorem can be derived from Theorem 1.1:
Theorem 1.2.
Suppose that $\Omega\subset\mathbb{R}^{3}$ satisfies Assumption A and that the kernel $B$ satisfies Assumption B with $\nu<0$. We consider a non-negative continuous function $f_{0}$ on $\overline{\Omega}\times\mathbb{R}^{3}$. Let $f(t,x,v)$ be a continuous mild solution of (1.1)–(1.3) on $0,T)\times\overline{\Omega}\times\mathbb{R}^{3}$, with initial condition $f_{0}$ for some $T>0$, and let $\alpha\in[0,1]$. We assume that $f(t,x,v)$ satisfies the following properties:
  1. (1)
    $M>0$;
  2. (2)
Then, the following lower bound holds: For any $\tau\in(0,T)$, there exist $\rho>0$ and $\theta>0$ depending on $M$, $C_{\Phi},c_{\Phi},\gamma,b_{0},\nu,E_{f}$ (and $L_{f,p}$ if $\gamma<0$ ), $\tau$, $\alpha$, $\Omega$ and on the modulus of continuity of $f_{0}$, such that
@@
For the non-cutoff case, the result is stated in Chapter 6. We note that the numbers $\rho$ and $\theta$ are computable thanks to the special zigzag which will be introduced in Proposition 2.5, through which we can guarantee that a certain amount and range of the lower bound can be generated through the propagation in Proposition 3.4 and derive a computable uniform lower bound (which depends on $\Omega$ since the construction of the zigzag is affected by the geometric properties of $\Omega$) as in Proposition 3.5.
The remainder of the article is organized as follows. In Section 2, we analyze the boundary of non-convex domain to derive some key properties. In Section 3, we focus on the derivation of a series of diluting initial lower bounds and the propagation effect of lower bounds. In Section 4 and Section 5, we provide a detailed proof of the Maxwellian lower bound on mild solution to the Boltzmann equation with non-fully specular ($0\leq\alpha<1$) and fully specular reflection boundary condition ($\alpha=1$), respectively. In Section 6, a weaker exponential lower bound of solutions to the Boltzmann equation for non-cutoff case is provided.